# Boundary Effects and Oxygen Deficiency-Driven Pattern Transitions in Algal Bioconvection

### v1

@@ 4 Discussion and Conclusions
Our findings are consistent with the recent study by Fragkopoulos et al. [31], who demonstrated that under anaerobic conditions, confined suspensions of photosynthetically active microorganisms exhibit reversibly switchable coherent flows. In their experiments, lowering the red-light intensity under anaerobic conditions—thereby reducing photosynthetic activity—led to a marked decrease in cell swimming speeds (from approximately 100,$\mu$m,s-1 to 40,$\mu$m,s-1) and initiated the emergence of spatially regular, temporally stationary bioconvective plumes within an otherwise stable suspension exhibiting an inverted sedimentation profile. While Fragkopoulos et al. systematically varied the light intensity to modulate metabolic activity, in our experiments the red-light intensity was held constant at approximately $5\times 10^{17}$ photons  m-2 s-1 throughout. Instead, oxygen availability was modulated purely through physical confinement. We propose that this confinement-induced oxygen depletion similarly drives a metabolic transition from aerobic to anaerobic conditions, ultimately leading to the bioconvection pattern transitions observed in our system. Thus, both work independently support the conclusion that metabolic activity and oxygen dynamics are key regulators of collective behavior in motile microbial suspensions and further emphasizes that geometric confinement alone, by limiting oxygen replenishment, can drive dramatic reorganizations of bioconvection patterns even in otherwise constant environmental conditions.
@@
Furthermore, an earlier study by Kage et al. [42] reported drastic reorganization of bioconvection patterns in confined CR suspensions under red-light illumination, which they primarily attributed to gyrotactic reorientation—namely, changes in the balance between gravitational and viscous torques experienced by the cells. In that study, the authors also observed that pattern wavelength shifts correlated with illumination intensity, noting that red light predominantly influenced the steady-state patterns after the transition, rather than affecting the initial onset or the pre-transition steady states. They proposed that these changes could be linked to alterations in cellular photosynthetic activity. In connection with these findings, our results—obtained under constant red-light illumination—highlight that confinement, and the associated oxygen depletion, likely inducing a metabolic shift from aerobic respiration to photosynthesis, serves as a critical driver of the observed pattern transitions. Notably, in our experiments, no pattern transitions were observed under open liquid–air interface conditions, nor in control experiments where the entire setup—including the top boundary—was fabricated from oxygen-permeable PDMS (Fig. S1). In contrast, striking pattern reorganizations consistently emerged under confined PMMA-based configurations with restricted oxygen replenishment, underscoring the critical role of oxygen availability in driving the observed transitions.
@@
To better understand the mechanisms underlying the pattern transitions observed in our experiments, we performed numerical simulations based on standard continuum models of bioconvection. These models, which couple the incompressible Navier–Stokes equations with cell conservation and gyrotactic reorientation [38, 43], have been widely used to describe bioconvection instabilities. In these frameworks, the onset of bioconvection is predicted when the Rayleigh number exceeds a critical threshold, with characteristic pattern wavelengths determined by the fastest-growing unstable modes. However, it is important to recognize the inherent limitations of these classical models. They typically assume constant swimming speeds and negligible cell–cell interactions under the dilute-suspension approximation. As noted previously [43, 15], in the nonlinear stages of bioconvection, local cell concentrations within dense, downwelling plumes can significantly exceed the initial mean value, challenging the validity of the dilute-suspension assumption and potentially altering hydrodynamic and collective behaviors. In our confined system, the observed pattern transitions—characterized by the temporary disappearance and subsequent reorganization into smaller-wavelength structures—cannot be explained solely by these classical bioconvection models. Even when our numerical simulations incorporated no-slip boundary conditions to emulate experimental confinement, the standard framework was only able to reproduce the initial formation and fragmentation of spiral-shaped cell accumulations into downwelling plumes; it failed to capture the striking reorganizations into new, stable patterns observed experimentally. This discrepancy highlights the need to extend standard bioconvection theories to account for additional factors, notably oxygen dynamics. Our experiments strongly suggest that oxygen depletion—and the associated transition from aerobic to anaerobic metabolism, which was shown to modulate microbial motility in Ref. [31]—plays a critical role in driving these transitions. Indeed, similar to the modeling of oxytactic bacterial suspensions [15, 44, 45, 46, 47], it may be necessary to incorporate an explicit oxygen concentration field $C_{oxy}(\bm{x},t)$, coupled to microbial motion. In these models, the average cell swimming velocity arises as a result of chemotaxis and is assumed to be directly proportional to the gradient of the oxygen concentration:
$W_{c}=\chi\nabla C_{oxy},$ (5)
where $\chi$ is the chemotactic sensitivity coefficient. Consequently, a further evolution equation for $C_{\text{oxy}}$ must be introduced, incorporating diffusion, advection by the fluid, and consumption by the microorganisms:
$\frac{\partial C_{\text{oxy}}}{\partial t}+\mathbf{U}.\nabla C_{\text{oxy}}=D_{\text{oxy}}\nabla^{2}C_{\text{oxy}}-\gamma c,$ (6)
where $D_{\text{oxy}}$ denotes the molecular diffusivity of oxygen (typically much larger than the effective diffusivity of the cells), $\gamma$ is the oxygen consumption rate per cell, and $c$ represents the local cell concentration. To fully capture the complex spatiotemporal dynamics and pattern transitions observed in confined active suspensions, future bioconvection models must explicitly incorporate oxygen transport and depletion. Moreover, the simple linear chemotactic response assumed in Eq. (5) may require refinement, as cellular swimming behavior could vary nonlinearly with oxygen concentration, particularly under conditions of severe depletion or strong spatial heterogeneity. Incorporating metabolically regulated motility responses will thus be critical for accurately describing the rich transitions uncovered in our experiments. Finally, potential crowding effects at elevated cell densities should also be considered to fully account for the intricate interplay between metabolic activity, confinement, and collective flow phenomena observed in our system.

### v3

@@ 4 Discussion and Conclusions
Our findings are consistent with the recent study by Fragkopoulos et al. 31, who demonstrated that under anaerobic conditions, confined suspensions of photosynthetically active microorganisms exhibit reversibly switchable coherent flows. In their experiments, lowering the red-light intensity under anaerobic conditions—thereby reducing photosynthetic activity—led to a marked decrease in cell swimming speeds (from approximately 100 $\mu$m s-1 to 40 $\mu$m s-1) and initiated the emergence of spatially regular, temporally stationary bioconvective plumes within an otherwise stable suspension exhibiting an inverted sedimentation profile. While Fragkopoulos et al. systematically varied the light intensity to modulate metabolic activity, in our experiments the red-light intensity was held constant at approximately $5\times 10^{17}$ photons  m-2 s-1 throughout. Instead, oxygen availability was modulated purely through physical confinement. We propose that confinement-induced oxygen depletion under low-light conditions triggers a metabolic shift from aerobic to anaerobic respiration, which in turn drives the observed transition of bioconvection patterns. Thus, both work independently support the conclusion that metabolic activity and oxygen dynamics are key regulators of collective behavior in motile microbial suspensions and further emphasizes that under low-light conditions geometric confinement alone, by limiting oxygen replenishment, can drive dramatic reorganizations of bioconvection patterns even in otherwise constant environmental conditions.
@@
Furthermore, an earlier study by Kage et al. 43 reported drastic reorganization of bioconvection patterns in confined CR suspensions under red-light illumination, which they primarily attributed to gyrotactic reorientation—namely, changes in the balance between gravitational and viscous torques experienced by the cells. In that study, the authors also observed that pattern wavelength shifts correlated with illumination intensity, noting that red light predominantly influenced the steady-state patterns after the transition, rather than affecting the initial onset or the pre-transition steady states. They proposed that these changes could be linked to alterations in cellular photosynthetic activity. In connection with these findings, our results—obtained under constant red-light illumination—highlight that confinement, and the associated oxygen depletion, likely inducing a shift from aerobic to anaerobic metabolism, serves as a critical driver of the observed pattern transitions. Notably, in our experiments, no pattern transitions were observed under open liquid–air interface conditions, nor in control experiments where the entire setup—including the top boundary—was fabricated from oxygen-permeable PDMS (Fig. S1). In contrast, striking pattern reorganizations consistently emerged under confined PMMA-based configurations with restricted oxygen replenishment, underscoring the critical role of oxygen availability in driving the observed transitions.
@@
To better understand the mechanisms underlying the pattern transitions observed in our experiments, we performed numerical simulations based on standard continuum models of bioconvection. These models, which couple the incompressible Navier–Stokes equations with cell conservation and gyrotactic reorientation 38, 44, have been widely used to describe bioconvection instabilities. In these frameworks, the onset of bioconvection is predicted when the Rayleigh number exceeds a critical threshold, with characteristic pattern wavelengths determined by the fastest-growing unstable modes. However, it is important to recognize the inherent limitations of these classical models. They typically assume constant swimming speeds and negligible cell–cell interactions under the dilute-suspension approximation. As noted previously 44, 15, in the nonlinear stages of bioconvection, local cell concentrations within dense, downwelling plumes can significantly exceed the initial mean value, challenging the validity of the dilute-suspension assumption and potentially altering hydrodynamic and collective behaviors. In our confined system, the observed pattern transitions—characterized by the temporary disappearance and subsequent reorganization into smaller-wavelength structures—cannot be explained solely by these classical bioconvection models. Even when our numerical simulations incorporated no-slip boundary conditions to emulate experimental confinement, the standard framework was only able to reproduce the initial formation and fragmentation of spiral-shaped cell accumulations into downwelling plumes; it failed to capture the striking reorganizations into new, stable patterns observed experimentally. This discrepancy highlights the need to extend standard bioconvection theories to account for additional factors, notably oxygen dynamics. Our experiments strongly suggest that oxygen depletion—and the associated transition from aerobic to anaerobic metabolism, which was shown to modulate microbial motility in Ref. 31—plays a critical role in driving these transitions.
Inspired by models of oxytactic bacterial suspensions and photosynthetic microbes (e.g., Refs. 15, 45, 46, 47, 48, 32), we introduce an explicit oxygen concentration field $C_{\mathrm{oxy}}(\mathbf{x},t)$ and couple it to the bioconvection flow $\mathbf{U}(\mathbf{x},t)$ and local cell concentration $c(\mathbf{x},t)$ via the reaction–diffusion–advection equation:
$\frac{\partial C_{\mathrm{oxy}}}{\partial t}+\mathbf{U}\cdot\nabla C_{\mathrm{oxy}}=D_{\mathrm{oxy}}\nabla^{2}C_{\mathrm{oxy}}-\gamma\,c\,C_{\mathrm{oxy}},$ (5)
where $D_{\mathrm{oxy}}=2000\,\mu\mathrm{m}^{2}\,\mathrm{s}^{-1}$ 49 is the molecular diffusivity of oxygen in water and $\gamma=10^{-5}\,\mathrm{mm}^{3}\,\mathrm{s}^{-1}$ is the assumed per-cell oxygen consumption rate. We then modulate the swimming speed $W_{c}$ based on the local oxygen concentration, increasing $W_{c}$ with oxygen levels via a sigmoidal function 32:
$W_{c}(C_{\mathrm{oxy}})=W_{\min}+\frac{W_{\max}-W_{\min}}{2}\Bigl{[}1+\tanh\!\bigl{(}\tfrac{C_{\mathrm{oxy}}-C_{\mathrm{typ}}}{\beta}\bigr{)}\Bigr{]},$ (6)
with parameters
$W_{\min}=56\,\mu\mathrm{m\,s}^{-1},\quad W_{\max}=112\,\mu\mathrm{m\,s}^{-1},$
and
$C_{\mathrm{sat}}=3\times 10^{-4}\,\mathrm{\mu mol\,mm}^{-3},\quad C_{\mathrm{typ}}=0.14\,C_{\mathrm{sat}},\quad\beta=0.11\,C_{\mathrm{sat}}.$
Here, $C_{\mathrm{sat}}$ denotes the saturation concentration of oxygen in water, and $C_{\mathrm{typ}}$ is a representative (typical) oxygen concentration. Simulations assume an initially uniform saturation, $C_{\mathrm{oxy}}(\mathbf{x},0)=C_{\mathrm{sat}}$.
Fig.  8: Top-view snapshots at $z=3\ \mathrm{mm}$ from 3D simulations of Eqs. (1)–(6) illustrate an anti-correlation between cell and oxygen concentrations; see Video 15 for the full dynamic evolution.
Figure 8 and Video 15 show the results of our simulations of Eqs. (1)–(6) using the parameters listed above, while all other parameters remain as in our previous simulations. We imposed a no-flux boundary condition for oxygen concentration and a no-slip condition for velocity on all domain boundaries . As expected, cell and oxygen concentrations are anti-correlated: regions with higher cell density exhibit lower oxygen levels. As oxygen levels fall, swimming speed declines, ultimately reaching the minimum value of $56\,\mu\mathrm{m}\,\mathrm{s}^{-1}$ prescribed by Eq. (6). However, our simulations indicate that oxygen dynamics alone cannot reproduce the dramatic phase transitions observed in our experiments, underscoring the necessity of more sophisticated models to accurately capture these complex transitions. Moreover, potential crowding effects at elevated cell densities should also be considered to fully account for the intricate interplay between metabolic activity, confinement, and collective flow phenomena observed in our system.
@@ Acknowledgements
Notes and references
  • Goldstein 2015 R. E. Goldstein, Annual review of fluid mechanics, 2015, 47, 343–375.
  • Brumley et al. 2014 D. R. Brumley, K. Y. Wan, M. Polin and R. E. Goldstein, elife, 2014, 3, e02750.
  • Wan and Goldstein 2016 K. Y. Wan and R. E. Goldstein, Proceedings of the National Academy of Sciences, 2016, 113, E2784–E2793.
  • Drescher et al. 2011 K. Drescher, J. Dunkel, L. H. Cisneros, S. Ganguly and R. E. Goldstein, Proceedings of the National Academy of Sciences, 2011, 108, 10940–10945.
  • Polin et al. 2009 M. Polin, I. Tuval, K. Drescher, J. P. Gollub and R. E. Goldstein, Science, 2009, 325, 487–490.
  • Ahmad et al. 2021 R. Ahmad, C. Kleineberg, V. Nasirimarekani, Y.-J. Su, S. Goli Pozveh, A. Bae, K. Sundmacher, E. Bodenschatz, I. Guido, T. Vidakovic-Koch and A. Gholami, ACS synthetic biology, 2021, 10, 1490–1504.
  • Mojiri et al. 2021 S. Mojiri, S. Isbaner, S. Mühle, H. Jang, A. J. Bae, I. Gregor, A. Gholami and J. Enderlein, Biomedical Optics Express, 2021, 12, 3169–3180.
  • Pedley and Kessler 1987 T. J. Pedley and J. Kessler, Proceedings of the Royal Society of London. Series B. Biological Sciences, 1987, 231, 47–70.
  • Pedley and Kessler 1992 T. Pedley and J. Kessler, Annual Review of Fluid Mechanics, 1992, 24, 313–358.
  • Klindt and Friedrich 2015 G. S. Klindt and B. M. Friedrich, Physical Review E, 2015, 92, 063019.
  • Lauga 2016 E. Lauga, Annual Review of Fluid Mechanics, 2016, 48, 105–130.
  • Kessler 1985 J. O. Kessler, Nature, 1985, 313, 218–220.
  • Bees and Hill 1998 M. A. Bees and N. Hill, Physics of Fluids, 1998, 10, 1864–1881.
  • Bees 2020 M. A. Bees, Annual Review of Fluid Mechanics, 2020, 52, 449–476.
  • Hill and Pedley 2005 N. Hill and T. Pedley, Fluid Dynamics Research, 2005, 37, 1.
  • Pedley and Kessler 1990 T. Pedley and J. O. Kessler, Journal of fluid mechanics, 1990, 212, 155–182.
  • Kessler 1985 J. O. Kessler, Contemporary Physics, 1985, 26, 147–166.
  • Bees and Hill 1997 M. A. Bees and N. Hill, Journal of experimental biology, 1997, 200, 1515–1526.
  • Williams and Bees 2011 C. R. Williams and M. A. Bees, Journal of Experimental Biology, 2011, 214, 2398–2408.
  • Javadi et al. 2020 A. Javadi, J. Arrieta, I. Tuval and M. Polin, Philosophical Transactions of the Royal Society A, 2020, 378, 20190523.
  • Dervaux et al. 2017 J. Dervaux, M. Capellazzi Resta and P. Brunet, Nature Physics, 2017, 13, 306–312.
  • Ramamonjy et al. 2023 A. Ramamonjy, P. Brunet and J. Dervaux, Journal of Fluid Mechanics, 2023, 971, A29.
  • Kage et al. 2020 A. Kage, T. Omori, K. Kikuchi and T. Ishikawa, Journal of Experimental Biology, 2020, 223, jeb205989.
  • Kessler 1986 J. O. Kessler, Journal of Fluid Mechanics, 1986, 173, 191–205.
  • Childress et al. 1975 S. Childress, M. Levandowsky and E. Spiegel, Journal of Fluid Mechanics, 1975, 69, 591–613.
  • Czirók et al. 2000 A. Czirók, I. M. Jánosi and J. O. Kessler, Journal of Experimental Biology, 2000, 203, 3345–3354.
  • Czirók and Vicsek 2000 A. Czirók and T. Vicsek, Physica A: Statistical Mechanics and its Applications, 2000, 281, 17–29.
  • Bearon et al. 2011 R. N. Bearon, A. Hazel and G. Thorn, Journal of fluid mechanics, 2011, 680, 602–635.
  • Bearon and Hazel 2015 R. N. Bearon and A. L. Hazel, Journal of Fluid Mechanics, 2015, 771, R3.
  • Karimi and Paul 2013 A. Karimi and M. R. Paul, Physical Review E—Statistical, Nonlinear, and Soft Matter Physics, 2013, 87, 053016.
  • Fragkopoulos et al. 2025 A. A. Fragkopoulos, F. Böhme, N. Drewes and O. Bäumchen, Proceedings of the National Academy of Sciences, 2025, 122, e2413340122.
  • Fragkopoulos et al. 2021 A. A. Fragkopoulos, J. Vachier, J. Frey, F.-M. Le Menn, M. G. Mazza, M. Wilczek, D. Zwicker and O. Bäumchen, Journal of the Royal Society Interface, 2021, 18, 20210553.
  • Harris et al. 2009 E. H. Harris, D. B. Stern and G. Witman, The chlamydomonas sourcebook, Elsevier San Diego, CA, 2009, vol. 1.
  • Catalan et al. 2025 R. E. Catalan, A. A. Fragkopoulos, A. Girot, M. Lorenz and O. Bäumchen, Nature Protocols, 2025, 1–26.
  • Berthold et al. 2008 P. Berthold, S. P. Tsunoda, O. P. Ernst, W. Mages, D. Gradmann and P. Hegemann, The Plant Cell, 2008, 20, 1665–1677.
  • Witman 1993 G. B. Witman, Trends in cell biology, 1993, 3, 403–408.
  • Rüffer and Nultsch 1990 U. Rüffer and W. Nultsch, Cell Motility and the Cytoskeleton, 1990, 15, 162–167.
  • Pedley and Kessler 1992 T. Pedley and J. O. Kessler, Annual Review of Fluid Mechanics, 1992, 24, 313–358.
  • Pedley et al. 1988 T. J. Pedley, N. Hill and J. O. Kessler, Journal of fluid mechanics, 1988, 195, 223–237.
  • Pedley 2010 T. Pedley, Experimental mechanics, 2010, 50, 1293–1301.
  • Pedley 2010 T. Pedley, Journal of fluid mechanics, 2010, 647, 335–359.
  • Aditi Simha and Ramaswamy 2002 R. Aditi Simha and S. Ramaswamy, Physical review letters, 2002, 89, 058101.
  • Kage et al. 2013 A. Kage, C. Hosoya, S. A. Baba and Y. Mogami, Journal of Experimental Biology, 2013, 216, 4557–4566.
  • Bees and Hill 1999 M. Bees and N. Hill, Journal of Mathematical Biology, 1999, 38, 135–168.
  • Keller and Segel 1971 E. F. Keller and L. A. Segel, Journal of theoretical biology, 1971, 30, 235–248.
  • Hopkins and Fauci 2002 M. M. Hopkins and L. J. Fauci, Journal of Fluid Mechanics, 2002, 455, 149–174.
  • Desai and Ardekani 2017 N. Desai and A. M. Ardekani, Soft Matter, 2017, 13, 6033–6050.
  • Hillesdon and Pedley 1996 A. Hillesdon and T. Pedley, Journal of Fluid Mechanics, 1996, 324, 223–259.
  • Ferrel and Himmelblau 1967 R. Ferrel and D. Himmelblau, J. Chem. Eng. Data, 1967, 12, 111–115.
5 Supplemental Figures
Fig.  S1: A control experiment was performed using a fully confined setup fabricated entirely from air-permeable PDMS. As illustrated in the corresponding space–time plots, no transition in the bioconvection patterns was observed even after two hours—markedly contrasting the transitions seen in confined PMMA-based configurations. This result underscores the critical role of oxygen availability in driving the observed pattern transitions; see Video 16.
Fig.  S2: Cell density distribution at the bottom of the setup ($z=0$) from 3D simulations of bioconvection patterns. Snapshots at $t=100$ s reveal that the cell concentration at the bottom is approximately half that near the surface (see Fig. 5), and the cell clusters remain more diffuse and unlike the near-surface structures, clusters at the bottom do not exhibit reorientation perpendicular to the spiral boundaries, highlighting the depth-dependent development of bioconvection patterns; see Videos 9 and 10 .
Fig.  S3: Top-view snapshots of cell density from the same 3D simulations presented in Fig. 5, but taken 1  mm above the spiral-shaped boundary at $z=5$ mm. This plane corresponds to the fluid-air interface in the left panels and to the rigid top wall boundary in the right panels. Either a zero-shear-stress condition (representing a free liquid-air interface) or a no-slip condition (representing a rigid top wall) is applied. As described in Fig. 5, CR cells swim upward (negative gravitaxis) and accumulate near the surface, forming a distinct spiral-shaped cell density pattern (A, D). This structure is dynamically unstable, fragmenting into smaller, high-density clusters that subsequently evolve into downwelling plumes; see also Fig. 7 and Video 17.